OpenStudy (anonymous):

Use this Polynomial Identity (x^2+y^2)^2=(x^2-y^2)^2+(2xy)^2 to help you create a list of 4 Pythagorean Triples. Here is a list of the ones I have already used to make 6 Pythagorean triples: (4,1), (2,1), (4,3), (7,3), (5,2), (3,1). I need 4 more, any help would be great!

OpenStudy (paxpolaris):

\[\left( x^2+y^2 \right)^2=\left( x^2-y^2 \right)^2+\left( 2xy \right)^2\]

OpenStudy (paxpolaris):

if we pick: \(x=2\), \(y=1\) .... we get \(5^2=3^2+4^2\) .... or the Pythagorean triplet (3,4,5)

OpenStudy (paxpolaris):

pick any 2 positive integers with \(x>y\) .... you'll keep getting Pythagorean triplets

OpenStudy (paxpolaris):

if we pick: \(x=3\) and \(y=2\), we get \(13^2=5^2+12^2 \to \large\left(5,12,13\right)\) if we pick: \(x=4\) and \(y=2\), we get \(20^2=12^2+16^2 \to \large(12,16,20)\) if we pick: \(x=5\) and \(y=3\), we get \(34^2=16^2+30^2 \to \large(16,30,34)\) if we pick: \(x=6\) and \(y=1\), we get \(37^2=35^2+12^2 \to \large(12,35,37)\)

OpenStudy (anonymous):

omg thank you so much! :)